OpenStudy (anonymous):

MEDAL ... Find a quadratic model for the set of values. (–2, –20), (0, –4), (4, –20)

OpenStudy (anonymous):

Given a general quadratic graph: y = ax^2 + bx + c Plugging in all 3 sets of values: (–2, –20), (0, –4), (4, –20) -20 = a(-2)^2 + b(-2) + c -20 = 4a - 2b + c ----- (1) -4 = a (0^2) + b(0) + c -4 = c --------(2) -20 = a(4)^2 + b (4) + c -20 = 16a + 4b + c ----- (3) Sub in c = -4, (1) becomes: -20 = 4a - 2b -4 ----> -16 = 4a - 2b ------(4) (3) becomes: -20 = 16a + 4b -4 ----> -16 =16a + 4b ------(5) Making 4a the subject: 4a = -16 + 2b Thus, subbing into (5) -16 = 4(4a) + 4b -16 = 4(-16+2b) + 4b -16 = -64 + 8b + 4b 48 = 12b b = 4 subbing b = 4 into (4) -16 = 4a -2(4) -16 = 4a -8 -8 = 4a a = -2 Thus, y = ax^2 + bx + c with a = -2, b = 4, c =-4